5. What mass of calcium chloride must be dissolved in water to renare 200 mL of a solution with concentration of 0.50 M? 6. W 5, 7, and 8 please!
Answer #1

5.

molarity = 0.5M

volume of solution = 200ml = 0.2L

no of moles of CaCl2 = molarity *volume in L

                                = 0.5*0.2   = 0.1moles

mass of CaCl2 = no of moles* gram molar mass

                       = 0.1*111

                        = 11.1g >>>answer

6.2H2S(g) + 3O2(g) -----------> 2SO2(g) + 2H2O(l)

B) 3

7.

N2 + 3H2 -----------> 2NH3

1 mole of N2 react with 3 moles of H2

2 moles of N2 react with = 3*2/1 = 6 moles of H2 is required

H2 is limiting reactant

3 moles of H2 react with exess of N2 to gives 2 moles of NH3

perccent yield   = actual yield *100/theoretical yield

                         = 1.5*100/2   = 75%

8. no of moles of H2O   = W/G.M.Wt

                                   = 2000/18   = 111.11moles

2H2O(l) ------------> 2H2(g) + O2

2 moles of H2O electrolysis to gives 1 mole of O2

111.11moles of H2O electrolysis to gives = 1*111.11/2   = 55.555moles of O2

n = 55.555moles

P = 1atm

T = 25+273 = 298K

PV = nRT

V   = nRT/P

      = 55.555*0.0821*298/1 = 1359L of O2 >>>>answer